OpenStudy (anonymous):

What are the values of x and y in this linear system? x + 2y = 12 2x + 3y = 8

OpenStudy (anonymous):

Solve one of the expressions for either the value of Y or X and substitute back into the other equation.

OpenStudy (anonymous):

So for example if I have the linear system, Equation 1 : X+3Y = 2 Equation 2 : 4Y + 2X = 1 I can choose either equation and solve for the X or Y so I will choose the first and solve for X X+3Y = 2 -----> X = 2-3Y Solved Equation 1 for X so I will substitute my new "X" into Equation 2 like so 4Y+2(2-3Y) = 1 And now all my terms are Y's and I can solve Y for an actually number and substitute that back into any of my 2 original equations and get a value for X. Using the same method solve your Linear system.